OpenStudy (anonymous):

A bakery sells cakes for $7 and pastries for $2. Phillip has $40. He would like to buy more cakes than pastries. If Phillip wants to buy the greatest number of cakes and pastries, which of the combinations below should he buy? A Cakes: 2 Pastries: 13 B Cakes: 4 Pastries: 2 C Cakes: 4 Pastries: 6 D Cakes: 5 Pastries: 2

OpenStudy (anonymous):

im thinking maybe D

OpenStudy (anonymous):

7*5=35+2**2=4 35+4=39 Correct answer is D

OpenStudy (anonymous):

thats what i thought